OpenStudy (anonymous):

Short Answer. Add the following polynomials: (-9x^2 + 7x + 9) + (-9x^3 + 8x^2 – 8)

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

\[-9x^3\] has no like term in the first polynomial \[8x^2-9x^2=-x^2\] \[9-8=1\] so you get \[-9x^3-x^2+1\]

OpenStudy (anonymous):

Thank you here's another one I need help on also